OpenStudy (smurfy14):

how do you write sqrt of -5 in form of bi where b is the real number?

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{-5}=\sqrt{-1*5}=\sqrt{-1}\sqrt{5}=i\sqrt{5}=\sqrt{5}i\] So \[\large \sqrt{-5}=\sqrt{5}i\]

OpenStudy (anonymous):

@smurfy note that the 'i" is outside of the radical, not inside!